A woman bought 65 lb. of fertilizer to spread over her rectangular backyard. if she spread it at a rate of 20 lb. to 1000 ft. and the length of the yard exceeds the width by 10 ft., what were the dimensions of the yard?
This is a word problem in a certain chapter in our book which deals all about Quadratic Equations. I hope you could help me out. Thanks.

Answer by stanbon(75887)

A woman bought 65 lb. of fertilizer to spread over her rectangular backyard. if she spread it at a rate of 20 lb. to 1000 ft. and the length of the yard exceeds the width by 10 ft., what were the dimensions of the yard?
-------
width = w ft
length = w+10 ft
-----
Area = w(w+10) = w^2 + 10w sq. ft.
--------------------------
Assuming that the 65 lb of fertilizer covers her yard,
area of yard = 65/(20 lb/1000 ft) = (65/20)*1000 = 3250 sq yds
-------------
Equation:
w^2 + 10w - 3250 = 0
w = 175.35 ft(width)
w+10 = 185.35 ft (length)
==============
Cheers,
Stan H.
---------------